# Koopman Operator Based Time-Delay Embeddings and State History Augmented LQR for Periodic Hybrid Systems: Bouncing Pendulum and Bipedal Walking

### v1

@@ Abstract
Time-delay embedding is a technique that uses snapshots of state history over time to build a linear state space model of a nonlinear smooth system. We demonstrate that periodic non-smooth or hybrid system can also be modeled as a linear state space system using this approach as long as its behavior is consistent in modes and timings. We extended time-delay embeddings to generate a linear model of two periodic hybrid systems—the bouncing pendulum and the simplest walker—with control inputs. This leads to a novel state history augmented linear quadratic regulator (LQR) which uses current and past state history for feedback control. Example code can be found at https://github.com/Chun-MingYang/koopman-timeDelay-lqr.git
@@ 1 Introduction
Periodic hybrid systems are systems that exhibit both continuous dynamics (like moving joints or limbs) and discrete events (like impacts, foot contacts, or switching control modes), and do so in a repeating, cyclic manner. Some examples robotic systems are systems that use legs to walk, run, climb; systems that use hands to brachiate, juggle, and swim; and manipulators that perform periodic pick and place. From a controls synthesis point of view, the hybrid nature of the dynamics are tackled by using a switching controller but it is often difficult to prove the stability of such a controller. In this paper, we use data-driven time delay embedding to identify a smooth model of periodic hybrid systems, which enables the use of provable stable and optimal controllers.
The traditional control approach is to choose two layers of control (Bhounsule et al., 2012; Plooij et al., 2014). First, create a nominal controller that achieves periodic solution. Second, linearize about the periodic solution and use a linear controller to achieve stability. This method does require switching the controller at every discontinuity and is limited to small perturbations around the periodic solution.
The Koopman operator provides a method to create a global linear model for a nonlinear system by projecting the dynamics in the function space (Koopman, 1931). The choice of the functions, also known as observables, is the key to getting the dynamics correctly represented. Given a choice of observables, a computationally efficient method known as Extended Dynamic Mode Decomposition (EDMD) can be used to compute the linear model. One can use neural network-based observables so that the discovery of the observables and the linear model can happen together (Lusch et al., 2018), but this makes EDMD nonlinear and thus, computationally challenging. So far, EDMD has been applied to create linear models of simple smooth systems such as quadcopters (Narayanan et al., 2023), underwater vehicles (Rahmani and Redkar, 2024), autonomous cars (Kim et al., 2001), and soft robot manipulators (Bruder et al., 2020).
This challenge becomes more pronounced when applied to hybrid systems due to their inherent transient features, such as discrete switching dynamics. However, a method known as Direct Encoding (DE) (based on a inner product of observables) was able to compute a global linearization of the hybrid dynamics using sufficiently high number of radial basis functions as observables (Asada, 2023). But the approximation broke down as the number of observables decreased indicating the sensitivity of the method to the choice and size of observables.
Time Delay embeddings use state history as Koopman observables to create a global linear model (Kamb et al., 2020; Brunton et al., 2017). Its basis is the Taken’s Embedding theorem (Takens, 1980), which states that the delay-embedded observables captures the structure of the original state space up to a diffeomorphism—meaning it preserves all topological and differential structures (such as smoothness, continuity, and dimensionality). The Time Delay Embedding have been used to model dynamical system with no control such as the van der Pols oscillator and the Lorenz system (Kamb et al., 2020).
@@
Taken’s Embedding Theorem, in its original form, does not apply to general non-smooth or hybrid systems. However, under certain specific conditions—such as when the switching behavior is periodic and consistent (i.e., the same sequence of modes and timing)—the delay observables can preserve the hybrid features by embedding the switching non-smoothness within the periodic structure (Navarrete and Viswanath, 2019). We take this approach to create a smooth model of a periodic hybrid system. Then we use a linear quadratic regulator to stabilize the system. The novelties of the work are as follows. (1) Time delay embedding is used to model a periodic hybrid systems while past works have only modeled smooth systems. (2) Time delay embedding is used to model a system with control inputs while past works have only modeled uncontrolled systems. (3) The development of the state history-augmented linear quadratic regulator (LQR), which uses current and past state for feedback control.
@@ 2.2 Time-delay Embedding with Koopman Observables
The time-delay embeddings offer another technique for obtaining a linear dynamical system in the function space. We collect the history of states and control inputs and pass these into the observable ${\bf g}({\bf z}_{i})$ and build the Hankel matrix, ${\bf H}_{i}\in{\mathbb{R}}^{p(N+1)\times p(M+1)}$
@@
where $\bf{H}^{\dagger}$ is the pseudo-inverse of ${\bf H}$ and $\mathbf{L}\in{\mathbb{R}}^{p(N+1)\times p(N+1)}$.
@@ 2.4 State History Augmented LQR Control
where ${\bf A}\in{\mathbb{R}}^{n(N+1)\times n(N+1)}$ and ${\bf B}\in{\mathbb{R}}^{n(N+1)\times m(N+1)}$ are the system matrices. Consider using controller ${\bf U}_{i}={\bf U}^{d}_{i}+\hat{\bf U}_{i}$, we can now use the linear quadratic regulator (LQR) to compute a gain ${\bf K}_{\mbox{\scriptsize LQR}}\in{\mathbb{R}}^{m(N+1)\times n(N+1)}$ such that the optimal correction control term $\hat{\bf U}_{i}=-{\bf K}_{\mbox{\scriptsize LQR}}{\bf\hat{X}_{i}}$ can be retrieved. It is important to note that unlike traditional LQR which uses the current state for feedback control, the controller uses the current state and past state history for feedback control. This ensures that the controller is retrospective and less reactive ensuring smoother control.
@@ 2.5 Application to Periodic Hybrid Systems
where $g,l,\lambda$ are gravity term, pendulum length, and damping constant respectively. The switching conditions $\mathcal{S}_{i}(\mathbf{x})=0$ are defined when pendulum passing through the given angles ±$\pm\theta^{*}$ with the discrete reset map $\mathcal{R}_{i}(\mathbf{x})$—an instantaneous opposite kick—defined as:
@@
$\mathcal{R}_{1}(\mathbf{x}):$$\begin{cases}\theta^{+}=-\theta^{*},\\[2.84526pt] \omega^{+}=\omega^{-}+\Delta\omega,\end{cases}$
@@
$\mathcal{R}_{2}(\mathbf{x}):$$\begin{cases}\theta^{+}=\theta^{*},\\[2.84526pt] \omega^{+}=\omega^{-}-\Delta\omega.\end{cases}$
where ,+-,+- , + are the notations representing the instance just before and just after the switch conditions; $\Delta\omega>0$ is the instantaneous kick modeled as instant augular velocity change. The limit cycle of the system can be obtained by considering $u=\lambda\omega$—which turns the system into an undamped pendulum that is naturally periodic—such that the periodic orbit, given $\mathbf{x}_{0}=[\theta_{0},\omega_{0}]^{T}$, can be obtained manually by selecting $\theta^{*}=\begin{Bmatrix}\theta(t_{s})|t_{s}\in[0,T]\end{Bmatrix}<max(\theta(0:T))$ and $\Delta\omega=2|\omega(t_{s})|$.
@@
[…]
@@
where ,+-,+- , + are the notations representing the instance just before and just after the foot strike event. The fixed points can be found by solving the following Poincaré analysis as:
@@ 3.2 Simplest Walker
Given the initial range $\theta_{0}\in[0.1,0.2],\phi_{0}\in[-0.4,-0.3]$ rad and $\dot{\theta}_{0}\in[-0.3,-0.2],\dot{\phi}_{0}\in[0.03,0.04]$ rad/s, along with the nominal controller $u_{1}=\lambda_{1}\dot{\theta},u_{2}=\lambda_{2}\dot{\phi}$ ($\lambda_{1}=\lambda_{2}=0.5$) to cancel the damping effect, the fixed points of the periodic walking limit cycle—whose Poincaré section is taken right after the foot-strike event—are found through Poincaré analysis to be $[\theta_{0},\phi_{0},\dot{\theta}_{0},\dot{\phi}_{0}]^{*}=[0.162,-0.325,-0.231,0.038]$ rad, such that the reference trajectory and norminal control profile, which also serve as the training dataset, are collected by simulating the model, initialed with Poincaré fixed points, for $8$ sec, as shown as dashed lines in Fig. 6.

### v2

@@ Abstract
Time-delay embedding is a technique that uses snapshots of state history over time to build a linear state space model of a nonlinear smooth system. We demonstrate that periodic non-smooth or hybrid system can also be modeled as a linear state space system using this approach as long as its behavior is consistent in modes and timings. We extend time-delay embeddings to generate a linear model of two periodic hybrid systems—the bouncing pendulum and the simplest walker—with control inputs. This leads to a state history augmented linear quadratic regulator (LQR) which uses current and past state history for feedback control. Example code can be found at https://github.com/Chun-MingYang/koopman-timeDelay-lqr.git
@@ 1 Introduction
Periodic hybrid systems are systems that exhibit both continuous dynamics and discrete events, and do so in a repeating, cyclic manner. Some examples of periodic hybrid robotic systems include legged robots that walk, run, or climb; robotic arms or hands that brachiate, juggle, or swim; and manipulators that perform periodic pick-and-place tasks. From a controls synthesis point of view, the hybrid nature of the dynamics are often tackled by using a switching controller but it is often difficult to guarantee the stability of such a controller. Another approach involves using a two-layer controller (Bhounsule et al., 2012; Plooij et al., 2014). The first layer designs a nominal controller to generate a periodic solution. The second layer then linearizes the dynamics around this periodic solution and applies a linear controller to ensure stability. However, this method requires switching controllers at each discontinuity and is effective only for small perturbations near the periodic trajectory. Recently, the development of the Koopman operator has offered a new perspective for analyzing and controlling the nonlinear systems.
The Koopman operator (Koopman, 1931) is a linear operator that evolves a nonlinear dynamics linearly in high-dimensional function space. While the original formulation was limited to ergodic systems, (Mezić, 2005) is the first work to extend the Koopman operator theory to general nonlinear systems, laying the foundation for later numerical methods—such as Dynamic Mode Decomposition (DMD) (Schmid, 2010; Williams et al., 2015; Proctor et al., 2016)—that approximate the Koopman operator in a data-driven fashion. The Koopman operator has since become a powerful tool in system identification and control. So far, the Koopman operator has been increasingly applied to robotic systems, including soft robot (Bruder et al., 2020), quadcopter (Narayanan et al., 2023), underwater robot (Rahmani and Redkar, 2024), and quadrupedal legged robot (Yang and Bhounsule, ; Yang and Bhounsule, 2025). However, most of the existing work has focused on continuous systems, with relatively few studies addressing hybrid systems.
The hybrid system poses greater challenges than the continuous system, in part because it is difficult to find suitable coordinates in high-dimensional space to capture its inherent transient features. The earliest work (Govindarajan et al., 2016) demonstrates that Koopman analysis can be extended to hybrid systems and shows that the hybrid dynamics can be evolved in the associated eigenspaces. Later, (Asada, 2023) proposed a practical method—using direct encoding with radial basis functions—to encode hybrid dynamics into a global linear operator model in the Hilbert space. Alternately, one can also use time-delay embedding to construct the high-dimensional space.
Time-delay embedding is a method used to lift the state space dimension by stacking a sequence of state history into a Hankel matrix based on measurements of generic observables—usually the state itself. The earliest work (Arbabi and Mezic, 2017) proved that one can find the Koopman operator by applying DMD to the time-delay embedded Hankel matrix, under the assumption of ergodic systems. Later, in the work of (Kamb et al., 2020), they established that projecting the time-delay coordinates onto an orthogonal basis yields a universal Koopman operator. (i.e., one invariant to the particular coordinate choice within the embedding.) This extension is grounded in Takens’ Embedding Theorem (Takens, 1980), which states that the time-delay embedding capture the structure of the original state space up to a diffeomorphism—meaning all topological and differential structures are preserved.
@@
Taken’s Embedding Theorem, in its original form, does not apply to general non-smooth or hybrid systems. However, under certain specific conditions—such as when the switching behavior is periodic and consistent (i.e., following the same sequence of modes and timing)—the delay observables can preserve the hybrid features by embedding the switching non-smoothness within the periodic structure (Navarrete and Viswanath, 2019). Building on this insight, we apply the time-delay embedding approach to construct a Koopman model for the periodic hybrid system with control inputs—a contrast to past works (Haggerty et al., 2023) that focused only on smooth systems. Once the model is obtained, we implement a linear quadratic regulator that leverages both current and past states for feedback control to stabilize the system.
@@ 2.2 Time-delay Embedding with Koopman Observables
The time-delay embeddings offer one technique for obtaining a linear dynamical system in the function space: We collect the history of states and control inputs and pass these into the observable ${\bf g}({\bf z}_{i})$ and build the Hankel matrix, ${\bf H}_{i}\in{\mathbb{R}}^{p(N+1)\times p(M+1)}$
@@
where $\bf{H}^{\dagger}$ is the pseudo-inverse of $\bf H$} and Koopman operator is $\mathbf{L}\in{\mathbb{R}}^{p(N+1)\times p(N+1)}$.
@@ 2.4 State History Augmented LQR Control
where ${\bf A}\in{\mathbb{R}}^{n(N+1)\times n(N+1)}$ and ${\bf B}\in{\mathbb{R}}^{n(N+1)\times m(N+1)}$ are the system matrices. Consider using controller ${\bf U}_{i}={\bf U}^{d}_{i}+\hat{\bf U}_{i}$, we can now use the linear quadratic regulator (LQR) to compute a gain ${\bf K}_{\mbox{\scriptsize LQR}}\in{\mathbb{R}}^{m(N+1)\times n(N+1)}$ such that the optimal correction control term $\hat{\bf U}_{i}=-{\bf K}_{\mbox{\scriptsize LQR}}{\bf\hat{X}_{i}}$ can be retrieved. It is important to note that unlike traditional LQR which uses the current state for feedback control, the controller uses the current state and past state history for feedback control. This ensures that the controller is retrospective and less reactive.
@@ 2.5 Application to Periodic Hybrid Systems
where $g,l,\lambda$ are gravity term, pendulum length, and damping constant respectively. The switching conditions $\mathcal{S}_{i}(\mathbf{x})=0$ are defined when pendulum passing through the given angles $\pm\theta^{*}$ with the discrete reset map $\mathcal{R}_{i}(\mathbf{x})$—an instantaneous opposite kick—defined as:
@@
[…]
@@
[…]
where ,+-,+ are the notations representing the instance just before and just after the switch conditions; $\Delta\omega>0$ is the instantaneous kick modeled as instant augular velocity change. The limit cycle of the system can be obtained by considering $u=\lambda\omega$—which turns the system into an undamped pendulum that is naturally periodic—such that the periodic orbit, given $\mathbf{x}_{0}=[\theta_{0},\omega_{0}]^{T}$, can be obtained manually by selecting $\theta^{*}=\begin{Bmatrix}\theta(t_{s})|t_{s}\in[0,T]\end{Bmatrix}<max(\theta(0:T))$ and $\Delta\omega=2|\omega(t_{s})|$.
@@
$\ddot{\theta}$ $=\sin(\theta-\gamma)-\lambda_{1}\dot{\theta}+u_{1}$ (20)
@@
where ,+-,+ are the notations representing the instance just before and just after the foot strike event. The fixed points can be found by solving the following Poincaré analysis as:
@@ 3.2 Simplest Walker
Given the initial range $\theta_{0}\in[0.1,0.2],\phi_{0}\in[-0.4,-0.3]$ rad and $\dot{\theta}_{0}\in[-0.3,-0.2],\dot{\phi}_{0}\in[0.03,0.04]$ rad/s, along with the nominal controller $u_{1}=\lambda_{1}\dot{\theta},u_{2}=\lambda_{2}\dot{\phi}$ ($\lambda_{1}=\lambda_{2}=0.5$) to cancel the damping effect, the fixed points of the periodic walking limit cycle—whose Poincaré section is taken right after the foot-strike event—are found through Poincaré analysis to be $\theta_{0},\phi_{0},\dot{\theta}_{0},\dot{\phi}_{0}]^{*}=[0.162,-0.325,-0.231,0.038]$ rad, such that the reference trajectory and norminal control profile, which also serve as the training dataset, are collected by simulating the model, initialed with Poincaré fixed points, for $8$ sec, as shown as dashed lines in Fig. 6.